# Vital Insight: Assisting Experts’ Context-Driven Sensemaking of Multi-modal Personal Tracking Data Using Visualization and Human-In-The-Loop LLM

### v3

@@ 9.7. Limitations
The study has several limitations worth noting. While we achieved data saturation, our small sample size limits the generalizability of our findings, particularly as it relates to other health and behavioral outcomes. Although recent research suggests small sample sizes are acceptable in interpretive studies (crabtree2024h, ), future work should involve larger populations and diverse scenarios to validate the findings. While researchers reviewed all the LLM-generated content shown in the two user studies to ensure basic correctness without making any changes, we conducted a preliminary quantitative evaluation of the LLM results. A more thorough evaluation of LLM results and capabilities can be conducted in future research, which is beyond the scope of this study. Another limitation arises from the fact that experts spent more time on tasks when LLM content was included compared to when it was not. In our qualitative analysis, we found that this was partly due to experts treating LLM content seriously in their sensemaking process, which led them to spend more time on it. Additionally, although we provided a tutorial video at the beginning, there was still a learning curve in using the information effectively. A more longitudinal study is needed to assess the long-term usage of the prototype in real-world settings and determine whether it can ultimately save experts time in sensemaking tasks. Although we conducted A/B testing using a baseline system without LLM, we did not compare our system with other dashboards due to the specificity of the task. As a result, we presented the survey results as descriptive measurements and highlighted the positive qualitative feedback from experts, which aligned with the survey findings.

### v4

@@ 9.7. Limitations
The study has several limitations worth noting. While we achieved data saturation, our small sample size limits the generalizability of our findings, particularly as it relates to other health and behavioral outcomes. Although recent research suggests small sample sizes are acceptable in interpretive studies (crabtree2024h, ), future work should involve larger populations and diverse scenarios to validate the findings. While researchers reviewed all the LLM-generated content shown in the two user studies to ensure basic correctness without making any changes, we conducted a preliminary quantitative evaluation of the LLM results. A more thorough evaluation of LLM results and capabilities can be conducted in future research, which is beyond the scope of this study. Another limitation arises from the fact that experts spent more time on tasks when LLM content was included compared to when it was not. In our qualitative analysis, we found that this was partly due to experts treating LLM content seriously in their sensemaking process, which led them to spend more time on it. Additionally, although we provided a tutorial video at the beginning, there was still a learning curve in using the information effectively. A more longitudinal study is needed to assess the long-term usage of the prototype in real-world settings and determine whether it can ultimately save experts’ time in sensemaking tasks. Although we conducted A/B testing using a baseline system without LLM, we did not compare our system with other dashboards due to the specificity of the task. As a result, we presented the survey results as descriptive measurements and highlighted the positive qualitative feedback from experts, which aligned with the survey findings.
@@ 10. Conclusion
Acknowledgments
This research is partially supported by the NIH National Institutes of Health, under award number P30DA029926, and the National Science Foundation, under award number CAREER 2442593, and 2112633, The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors. Any mention of specific companies or products does not imply any endorsement by the authors, by their employers, or by the sponsors.